# Wave Attenuation in Drifting Sea Ice: A Mechanistic Model for Observed Decay Profiles

### v1

@@ 3 Comparisons to Antarctic Measurements
In this section comparisons are made with measured amplitudes and derived attenuations in the Antarctic sea ice (Brouwer et al., 2022; Voermans et al., 2025). The data comprises quality controlled ICESat-2 satellite transects of wave height (total and across frequency components corresponding to periods $T=9,12,15,18\text{ s}$) across the MIZ throughout 2019. Note that transects seldom start at the ice edge and, in computing attenuation rates, transects are assumed to align with the wave direction (Voermans et al., 2025). Drift is not measured by ICESat-2.
To compare data to model results, which are derived for a monochromatic wave, we attribute all energy to the most energetic frequency component measured over the transect, normalising amplitude and attenuation rate by the first available measurement from the ice edge, $a_{V}^{1}$ and $\alpha^{1}_{V}$ respectively. The ice edge is taken to coincide be the start of each transect.
Drift and drag are used as fitting parameters, within the range of available Antarctic sea ice observations, i.e. $v=\mathcal{O}(10^{-1})\,\text{ m/s }$ and $C_{d}=\mathcal{O}(10^{-2})$, together with $\alpha$. The drift is assumed to be constant along the transect, justified by slow spatial variation over 10–100 km in Antarctic sea ice (Holland & Kwok, 2012). The water density is set $\rho=1025\text{ kg/m}^{3}$ and deep water dispersion relation used, namely $\Omega=\omega$.
@@ 3.1 Individual Transects
The two transects reported in Voermans et al. (2025) are analysed in detail. Transect I, taken on 26 December 2019 at $85^{\circ}\mathrm{W},\;69^{\circ}\mathrm{S}$, is 61.5 km long and waves have peak period 15 s. Transect II, taken on 24 May 2019 at $88^{\circ}\mathrm{W},\;68^{\circ}\mathrm{S}$, is longer ($x_{\text{MIZ}}=137.5\text{ km}$) and waves have a shorter peak period (T=T=12 s).
For Transect I the parameters $v=0.26\text{ m/s},C_{d}=6.0\times 10^{-3}$ and $\alpha=5.0\times 10^{-6}$ provide good qualitative and quantitative agreement in terms of amplitude decay (see Fig. 2a). Similar performances are achieved by the the non-drifting model, i.e. $v=0$, and the exponential decay. However, the model with drift outperforms the others reproducing the observed increase in effective attenuation rate along the transect (between $\hat{x}/x_{\text{MIZ}}\approx 0.2\text{ and }0.6$; see Fig. 2b). Note that for the chosen parameters, drift is stronger than the orbital velocity throughout the transect ($\lvert v\rvert/U_{0}\approx 1.2$), i.e. only the branch $x>\hat{x}^{*}$ of Eq. 15 is used. Remarkably, the model predicts an extinction location consistent with the observed wave-affected sea ice extend, i.e. at $x/x_{\text{MIZ}}\approx 1$.
@@
For Transect II the parameters $v=0.49\text{ m/s},C_{d}=2.8\times 10^{-3}$ and $\alpha=1.0\times 10^{-7}$ qualitatively reproduce the measured mild amplitude decay up to $\hat{x}/x_{\text{MIZ}}<0.6$ followed by a sharp decline (see Fig. 2c). Other models, i.e. the exponential and non-drifting model, fail to capture the sharp decline deeper into the sea ice. The predictive performance of the drifting model is clearly highlighted by the attenuation rate, shown in Fig. 2d, where steady attenuation rate near the ice edge followed by rapid growth for $x/x_{\text{MIZ}}>0.6$ are well captured qualitatively and quantitatively. A jump discontinuity appears at $x/x_{\text{MIZ}}\approx 0.1$, consistent with the choice of $v$ resulting in $\lvert v\rvert/U_{0}\approx 0.9$ and the presence of $\hat{x}^{*}$ within the domain. Unlike our formulation, all other models predict either a decreasing attenuation rate (no drift model) or constant attenuation (exponential model), as shown in Fig. 2d.
In summary, Transect I and II showcase two different regimes, in Transect I drag attenuation is weak ($\delta=0.6$) whereas in Transect II it is the dominant mechanism ($\delta=8.0$). Therefore, as expected, in Transect I deviations from exponential attenuation are weaker than in Transect II however, also in Transect II, inclusion of exponential attenuation through $\alpha$ provides significantly better predictions. When drag dominates, the extinction location $x_{\text{end}}$ occurs earlier and the spike in attenuation is more prominent.
@@ 3.2 Averaged Attenuation
Voermans et al. (2025) report average attenuation profiles across seasons and sectors of the Antarctic MIZ. To reproduce the observed variability, we drive our model, i.e. Eq. 15, with $N_{0}=3\times 10^{5}$ random realisations with Gaussian distributed drift velocity (with mean $\mu_{v}$ and standard deviation $\sigma_{v}$), constant throughout the transect. The simulations replicate the variability of drift around Antarctica (Kwok et al., 2017). The attenuation rate is computed for the wave component T=12sT=12\ $\mathrm${s} (angular frequency $\Omega=0.52\,\mathrm{s}^{-1}$ and wavelength $\lambda=225$ m) and we set the wave amplitude at the sea ice edge to $a_{0}=1\text{m}$. The other parameters are $C_{d}=0.02$ and $\alpha=7.0\times 10^{-6}\ \mathrm{m}^{-1}$, to match the averaged measured attenuation rate at the sea ice edge. For the comparison, only $a\geq 0.05\text{m}$ are averaged, in analogy with the methodology used by Voermans et al. (2025) to exclude data with high uncertainty.
For drift $\mu_{v}=0.22$ m/s ($\mu_{v}/U_{0}=0.42$) and variance $\sigma_{v}=0.03$ m/s (Fig. 3a) the attenuation only slightly increases up to $x/\lambda\approx 550$, consistent with the fact that simulations are in weak drag regime ($\delta=0.42$). From $x/\lambda\approx 550$ to $x/\lambda\approx 700$ the increase in more noticeable, up to $\approx 3$ times the attenuation rate at the sea ice edge, in qualitative agreement with simulations for Transect A (cf. Fig. 2b), before stabilising farther in the sea ice domain. The spread of individual profiles, shaded in grey, also grows when attenuation rate increases. Note that the increase in attenuation corresponds to a reduction in available measurements of amplitude over threshold (shown in red) and, similarly, individual amplitude profiles reaching the extinction location (distribution shaded in blue).
Figure 3: Averaged attenuation rates (in black; for negative velocity in green) and interquartile range (shaded in grey). On the right axis, number of measurements (in red) and distribution of the extinction location (shaded in blue; not to scale). Insets show the drift velocity distribution.
@@
The increase in drift ($\mu_{v}=0.29$ m/s, $\mu_{v}/U_{0}=0.55$ and $\sigma_{v}=0.03$ m/s; Fig. 3c), results in higher $\delta=0.55$, still in weak drag regime, and moves the increase in attenuation rate closer to the sea ice edge, from $x/\lambda\approx 450$ to $x/\lambda\approx 550$. A slightly higher attenuation is also achieved compared to the previous cases, up to $\alpha_{\textrm{eff}}\approx 6\times 10^{-4}$, after which decreases slightly. The number of observations decreases sharply due to individual transects reaching their extinction location (cf. Fig. 1b), and only less than 0.1% remain after 800 wavelengths, also explaining the larger variance in effective attenuation.
In Fig. 3 the mean attenuation rates for negative mean velocity are also reported in dashed green line. The direction of the drift has only marginal effect on the effective attenuation profile, i.e. the growth of attenuation rate is deeper into the sea ice (by $\approx 10$ wavelengths) but does not affect its magnitude. This is expected for oceanic conditions, as $\alpha\approx\alpha_{\textrm{exp}}$ and $\Gamma\approx C_{d}/(gc_{g}2\pi)$, hence the dependence of direction of the drift is negligible for operational purposes.
Fig. 3c is remarkably similar to the measured averaged attenuation in Voermans et al. (2025), i.e. a linear increase followed by slight decline, also matching the range of observed attenuation rates. Unlike in the measurements, in the model all individual transects are available from the edge but the decrease of available observations in Voermans et al. (2025) deeper into the sea ice is also captured. It can be argued that the presence of drift limits the extent of the wave-affected area, and its extent in our model ($\approx 185$ km; Fig. 3d) is consistent to the one in observations (Brouwer et al., 2022) and model simulations (Day et al., 2024).
@@ 4 Conclusions
We derived a mechanistic model for wave attenuation that accounts for both phenomenological exponential attenuation mechanisms and dissipation due to drag at the ocean-sea ice interface, as well as the effect of sea ice drift, therefore combining and extending previous models (cf. Shen & Squire, 1998; Kohout et al., 2011; Herman et al., 2019). We identified the parameter $\delta$ which governs attenuation dominated by drift-induced drag mechanisms ($\delta>1$) and exponential mechanisms ($\delta<1$) . Analytical solutions reveal that in drag dominated regime, the amplitude eventually dies-off, unlike prediction based on exponential attenuation, and the attenuation profile spikes close to the extinction location.
@@
The model agrees qualitatively and quantitatively well with pan-Antarctic measurements of attenuation rate recently reported by Voermans et al. (2025) when sea ice is allowed to drift slowly ($\mu_{v}/U_{0}=0.55$), even for small $\delta$, i.e. $\delta=0.55$. In particular, drift is responsible for the increase in wave attenuation rate deeper into the sea ice and, to a certain extent, regulates the extent of the wave-affected sea ice in quantitative agreement to observations (Brouwer et al., 2022) and model simulations (Day et al., 2024).
@@
The quantitative agreement against individual transects is reasonable with limitations. Differences with the measured data can be attributed to experimental uncertainty, e.g. negative attenuation rates and misalignment with wave direction, and current model assumptions, i.e. no interaction between wave components, no presence of other forcing mechanisms, heterogeneous drift along the transect and changes in ice type and morphology.
@@
Ultimately the model provides a solid theoretical foundation to better interpret observations of wave attenuation in drifting sea ice as found around Antarctica, and prompts for new measurements for further validation. Moreover, while only drag has been studied explicitly, the approach can be extended to better specify other wave attenuation mechanisms for integration in coupled wave-sea ice models.

### v2

@@ 3 Comparisons to Antarctic Measurements
In this section comparisons are made with measured amplitudes and derived attenuations in the Antarctic sea ice (Brouwer et al., 2022; Voermans et al., 2025). The observations comprises quality controlled ICESat-2 satellite transects of wave height (both the total wave heights and the spectral wave heights corresponding to frequency bands at periods $T=9,12,15,18\text{ s}$) across the MIZ throughout 2019. Note that transects seldom start at the ice edge (possibly because they fail quality control in intermediate sea ice concentrations) and, in computing attenuation rates, transects are assumed to align with the wave direction (Voermans et al., 2025). Drift is not measured by ICESat-2.
To compare observations to model results, which are derived for a monochromatic wave, we attribute all energy to the most energetic frequency component measured over the transect, normalising amplitude and attenuation rate by the first available measurement from the ice edge, $a_{V}^{1}$ and $\alpha^{1}_{V}$. The ice edge is taken to coincide be the start of each transect. To allow for a straightforward comparison of the model with the observations we use $x$ in place of $\hat{x}$ noting that their difference is negligible for small drift velocities. Moreover, in agreement with the observations, transects are normalised by MIZ extent $x_{\text{MIZ}}$. In this regard is important to note that MIZ extent is reported independently from the wave amplitude, i.e. low amplitude readings were excluded due to quality control (Voermans et al., 2025).
Drift and drag are used as fitting parameters, within the range of available Antarctic sea ice observations, i.e. $v=\mathcal{O}(10^{-1})\,\text{ m/s }$ and $C_{d}=\mathcal{O}(10^{-2})$, together with $\alpha$. While least squares optimisations was attempted, the noise in the observational data prevented reasonable fitting throughout the transect. For illustrative purposes, we performed manual tuning for each transect of the parameters within ranges of characteristic values for the MIZ. The drift is assumed to be constant along the transect, justified by slow spatial variation over 10–100 km in Antarctic sea ice (Holland & Kwok, 2012). The water density is set $\rho=1025\text{ kg/m}^{3}$ and deep water dispersion relation used, namely $\Omega=\omega$.
@@ 3.1 Individual Transects
The two transects reported in Voermans et al. (2025) are analysed in detail. Transect I, taken on 26 December 2019 at $85^{\circ}\mathrm{W},\;69^{\circ}\mathrm{S}$, is 61.5 km long and waves have peak period 15 s with amplitude $a^{1}_{V}=0.45\text{ m and attenuation},\alpha^{1}_{\mathrm{V}}=1.42\times 10^{-5}\,\mathrm{m}^{-1}$. Transect II, taken on 24 May 2019 at $88^{\circ}\mathrm{W},\;68^{\circ}\mathrm{S}$, is longer ($x_{\text{MIZ}}=137.5\text{ km}$) and waves have a shorter peak period (T=T=12 s) with amplitude $a^{1}_{V}=0.55\text{ m and attenuation},\alpha^{1}_{\mathrm{V}}=2.43\times 10^{-6}\,\mathrm{m}^{-1}$.
For Transect I the parameters $v=0.26\text{ m/s},C_{d}=6.0\times 10^{-3}$ and $\alpha=5.0\times 10^{-6}$ provide good qualitative and quantitative agreement in terms of amplitude decay (see Fig. 2a). Similar performances are achieved by the the non-drifting model, i.e. $v=0$, and the exponential decay. However, the model with drift outperforms the others reproducing the observed increase in effective attenuation rate along the transect (between $x/x_{\text{MIZ}}\approx 0.2\text{ and }0.6$; see Fig. 2b). Note that for the chosen parameters, drift is stronger than the orbital velocity throughout the transect ($\lvert v\rvert/U_{0}\approx 1.2$), i.e. only the branch $x>\hat{x}^{*}$ of Eq. 15 is used. Remarkably, the model predicts an extinction location consistent with the observed wave-affected sea ice extend, i.e. at $x/x_{\text{MIZ}}\approx 1$.
@@
For Transect II the parameters $v=0.49\text{ m/s},C_{d}=2.8\times 10^{-3}$ and $\alpha=1.0\times 10^{-7}$ qualitatively reproduce the measured mild amplitude decay up to $x/x_{\text{MIZ}}<0.6$ followed by a sharp decline (see Fig. 2c). Other models, i.e. the exponential and non-drifting model better reproduce the attenuation near to the ice edge, but fail to capture the sharp decline deeper into the sea ice. The predictive performance of the drifting model is clearly highlighted by the attenuation rate, shown in Fig. 2d, where steady attenuation rate near the ice edge followed by rapid growth for $x/x_{\text{MIZ}}>0.6$ are well captured qualitatively and quantitatively. Note that the observed attenuation rate fluctuates resulting in negative values close to the edge and again at $x/x_{\text{MIZ}}\approx 0.4$, and models can not capture this behaviour. A jump discontinuity appears at $x/x_{\text{MIZ}}\approx 0.1$, consistent with the choice of $v$ resulting in $\lvert v\rvert/U_{0}\approx 0.9$ and the presence of $\hat{x}^{*}$ within the domain. Unlike our formulation, all other models predict either a decreasing attenuation rate (no drift model) or constant attenuation (exponential model), as shown in Fig. 2d.
In summary, Transect I and II showcase two different regimes, in Transect I drag attenuation is weak ($\delta=0.6$) whereas in Transect II it is the dominant mechanism ($\delta=8.0$). Therefore, as expected, in Transect I deviations from exponential attenuation are weaker than in Transect II however, also in Transect II, inclusion of exponential attenuation through $\alpha$ provides significantly improved predictions. When drag dominates, the extinction location $x_{\text{end}}$ occurs earlier and the spike in attenuation is more prominent.
Our model is developed for homogeneous sea ice conditions, implying constant $C_{d}$ along the transect. While this hypothesis is more likely to hold in the winter season (Transect II) its applicability to Transect I is more debatable. Moreover, direct comparison with individual observations is made difficult by inherent measurement uncertainty, including misalignenment of the transect with the wave direction Voermans et al. (2025), lack of information on the local sea ice properties, and possible presence of other forcing mechanisms in the wave transport equation, e.g. wind forcing. Nevertheless, we highlight that our simple model captures the observed feature of wave attenuation in the MIZ.
@@ 3.2 Averaged Attenuation
Voermans et al. (2025) report average attenuation profiles across seasons and sectors of the Antarctic MIZ. To reproduce the observed variability, we drive our model, i.e. Eq. 15, with $N_{0}=3\times 10^{5}$ random realisations with Gaussian distributed drift velocity (with mean $\mu_{v}$ and standard deviation $\sigma_{v}$), constant throughout the transect. Each realisation corresponds to an individual transect and $N$ denotes the number of transects at any given location according to our model, the decrease is due to individual transects reaching the extinction location. The simulations replicate the variability of drift around Antarctica (Kwok et al., 2017). The attenuation rate is computed for the wave component T=12sT=12\ $\mathrm${s} (angular frequency $\Omega=0.52\,\mathrm{s}^{-1}$ and wavelength $\lambda=225$ m) and we set the wave amplitude at the sea ice edge to $a_{0}=1\text{m}$. The other parameters are $C_{d}=0.02$ and $\alpha=7.0\times 10^{-6}\ \mathrm{m}^{-1}$, to match the averaged measured attenuation rate at the sea ice edge. While the drag coefficient is higher than the one chosen for the individual transects in the previous section, its value is within the range of observed values for the MIZ (McPhee, 1979). Moreover, the chosen value allows to quantitatively match the observed attenuation rate close to the sea ice edge as reported in Voermans et al. (2025). Smaller values of drag would give rise to a similar qualitative behaviour. For the comparison, only $a\geq 0.05\text{m}$ are averaged, in accordance with the methodology used by Voermans et al. (2025) to exclude observations with high uncertainty.
For drift $\mu_{v}=0.22$ m/s ($\mu_{v}/U_{0}=0.42$) and standard deviation $\sigma_{v}=0.03$ m/s (Fig. 3a) the attenuation only slightly increases up to $x/\lambda\approx 550$, consistent with the fact that simulations are in weak drag regime ($\delta=0.42$). From $x/\lambda\approx 550$ to $x/\lambda\approx 700$ the increase is more noticeable, up to $\approx 3$ times the attenuation rate at the sea ice edge, in qualitative agreement with simulations for Transect I (cf. Fig. 2b), before stabilising farther in the sea ice domain. The spread of individual profiles, shaded in grey, also grows when attenuation rate increases. Note that the increase in attenuation corresponds to a reduction in available measurements of amplitude over threshold (shown in red) and, similarly, individual amplitude profiles reaching the extinction location (distribution shaded in blue).
Figure 3: Averaged attenuation rates (in black; for negative velocity in green) and interquartile range (shaded in grey). On the right axis, number of measurements (in red) and distribution of the extinction location (shaded in blue; not to scale). Insets show the corresponding drift velocity distribution: (a) $\mu_{v}=0.22$  m s-1 $\sigma_{v}=0.03$  m s-1; (b) $\mu_{v}=0.22$  m s-1 $\sigma_{v}=0.01$  m s-1; (c) $\mu_{v}=0.29$  m s-1 $\sigma_{v}=0.03$  m s-1.
@@
The increase in drift ($\mu_{v}=0.29$ m/s, $\mu_{v}/U_{0}=0.55$ and $\sigma_{v}=0.03$ m/s; Fig. 3c), results in higher $\delta=0.55$, still in weak drag regime, and moves the increase in attenuation rate closer to the sea ice edge, from $x/\lambda\approx 450$ to $x/\lambda\approx 550$. A slightly higher attenuation is also achieved compared to the previous cases, up to $\alpha_{\textrm{eff}}\approx 6\times 10^{-4}$, after which decreases slightly. For smaller drag coefficients results the increase in attenuation occurs further into the ice field. The number of observations decreases sharply due to individual transects reaching their extinction location (cf. Fig. 1b), and only less than 0.1% remain after 800 wavelengths, also explaining the larger variance in effective attenuation.
In Fig. 3 the mean attenuation rates for negative mean velocity are also reported in dashed green line. The direction of the drift has only marginal effect on the effective attenuation profile, i.e. the growth of attenuation rate is deeper into the sea ice (by $\approx 10$ wavelengths) but does not affect its magnitude, with similar effect on available measurements and extinction location distributions. This is expected for oceanic conditions, as $\alpha\approx\alpha_{\textrm{exp}}$ and $\Gamma\approx C_{d}/(gc_{g}2\pi)$, hence the dependence of direction of the drift is negligible for operational purposes.
Fig. 3c is remarkably similar to the measured averaged attenuation in Voermans et al. (2025), i.e. a linear increase followed by slight decline, also matching the range of observed attenuation rates. Unlike in the measurements, in the model all individual transects are available from the edge but the decrease of available observations in Voermans et al. (2025) deeper into the sea ice is also captured. It can be argued that the presence of drift limits the extent of the wave-affected area, and its extent in our model ($\approx 185$ km; Fig. 3c) is consistent to the MIZ width of the order of 200 km as reportedin observations (Brouwer et al., 2022) and model simulations (Day et al., 2024).
@@ 4 Conclusions
We derived a mechanistic model for wave attenuation that accounts for both phenomenological exponential attenuation mechanisms and dissipation due to drag at the ocean-sea ice interface, as well as the effect of sea ice drift, therefore combining and extending previous models in which drift was excluded (cf. Shen & Squire, 1998; Kohout et al., 2011; Herman et al., 2019). We identified the parameter $\delta$ which governs attenuation dominated by drift-induced drag mechanisms ($\delta>1$) and exponential mechanisms ($\delta<1$). Analytical solutions reveal that in drift-induced drag dominated regime, the amplitude eventually dies off, unlike prediction based on exponential attenuation, and the attenuation profile spikes close to the extinction location.
@@
The model agrees well, both qualitatively and quantitatively with pan-Antarctic measurements of attenuation rate recently reported by Voermans et al. (2025) when sea ice is allowed to drift slowly ($\mu_{v}/U_{0}=0.55$), even for small $\delta$, i.e. $\delta=0.55$. In particular, drift is responsible for the increase in wave attenuation rate deeper into the sea ice and, to a certain extent, regulates the width of the wave-affected sea ice which is in quantitative agreement to the MIZ extent (100–200 km) reported for observations (Brouwer et al., 2022) and model simulations (Day et al., 2024).
@@
The quantitative agreement against individual transects is reasonable with limitations. Differences with the measured observations can be attributed to experimental uncertainty and current model assumptions, i.e. no interaction between wave components, no presence of other forcing mechanisms, homogeneous drift and sea ice properties along the transect. Some of the limitations of the model could be sought by seeking numerical solutions for the governing equation (Eq. 5) and allowing for spatially heterogenous parameters. In this regard, our modelling framework allows for easy implementation into operational models and the biggest limitation, common to many wave attenuation models, is the lack of physically-informed sea ice parameters along the MIZ.
@@
Ultimately the model provides a solid theoretical foundation to better interpret observations of wave attenuation in drifting sea ice as found around Antarctica, and prompts new concurrent measurements of waves, drift, and sea ice properties for further validation. Moreover, while only drag has been studied explicitly, the approach can be extended to better specify other wave attenuation mechanisms for integration in emerging coupled ocean-wave-sea ice models as the ones currently being developed in most weather forecasting agencies (Kousal et al., 2026), e.g. ECMWF.
Appendix A Integral Approximation
The integral II in Eq.7:
$I=\int_{0}^{2\pi}|U\sin\phi-v|^{3}\,d\phi,$ (17)
is treated differently for $|v|\ll a\Omega$ (or equivalently $|v|\ll U$) and $|v|>a\Omega$ (or equivalently $|v|>U$).
A.1 Case $|v|\ll U$.
We expand the integrand as:
$|U\sin\phi-v|^{3}=|U\sin\phi|^{3}-3vU^{2}\sin^{2}\phi\,\mathrm{sgn}(\sin\phi)\sin\phi+3v^{2}U|\sin\phi|+O(v^{3}).$ (18)
The second term has zero contribution towards $I$. Therefore, we approximate the integral as a polynomial in $U$:
$I\approx\int_{0}^{2\pi}\left(U^{3}|\sin\phi|^{3}+3Uv^{2}|\sin\phi|\right)d\phi=\dfrac{8}{3}U^{3}+12Uv^{2}$ (19)
While strictly valid only for $|v|\ll U$ we use this approximation for $|v|\leq U$.
A.2 Case: $|v|>U$
The integrand only depends on the sign of $v$. Therefore it can be written as:
$|U\sin\phi-v|^{3}=\text{sign}(v)\,(v-U\sin\phi)^{3}.$ (20)
Note that some of the terms have zero contribution towards $I$ due to symmetry. We write the integral as a polynomial:
$I=\int_{0}^{2\pi}\left(|v|^{3}+3|v|U^{2}\sin^{2}\phi\right)d\phi$ (21)
$I=2\pi|v|^{3}+3\pi U^{2}|v|$ (22)
Appendix B Analytical Solutions
The integral approximation $I$ is piecewise, i.e. depends on the relationship between $U$ and $v$. As a result, different governing equations emerge for solutions in region A, corresponding to $a\Omega\geq|v|$ , and region B for $a\Omega<|v|$. We first solve the governing equation for $a$ in region A and then impose continuity on the amplitude to obtain the solution in region B.
B.1 Region A.
In Region A, defined as the region in which orbital velocity is greater than the drift ($a\Omega\geq|v|$), if it exists we impose the initial condition $a(\hat{x}=0)=a_{0}$.
The governing equation in its differential and integral form is:
$\frac{da_{A}}{d\hat{x}}=-\alpha a_{A}-\tfrac{8}{3}\Gamma\Omega^{3}a_{A}^{2}-12\Gamma\Omega|v|^{2}\equiv-f(a),$ (23)
$\int_{a_{0}}^{a_{A}}\frac{da^{\prime}}{\tfrac{8}{3}\Gamma\Omega^{3}a^{\prime 2}+\alpha a^{\prime}+12\Gamma\Omega|v|^{2}}=-\hat{x}$. (24)
The sign of the discriminant of $f$, denoted as $\Delta$, sets the form of solution.
The roots of $f$ are $r_{\pm}$.
The solutions are categorised according to $\Delta$:
$a_{A}(\hat{x})=\begin{cases}\displaystyle\frac{3\sqrt{\Delta}}{16\Gamma\Omega^{3}}\tan\Biggl(-\frac{\sqrt{\Delta}}{2}\hat{x}+\arctan\Bigl(\frac{16\Gamma\Omega^{3}a_{0}+3\alpha}{3\sqrt{\Delta}}\Bigr)\Biggr)-\frac{3\alpha}{16\Gamma\Omega^{3}},&\Delta>0,\\[6.45831pt] \displaystyle\frac{32\Gamma\Omega^{3}a_{0}-\left(16\alpha\Gamma\Omega^{3}a_{0}+3\alpha^{2}\right)\hat{x}}{32\Gamma\Omega^{3}+\left(32\Gamma\Omega^{3}a_{0}+6\alpha\right)\hat{x}},&\Delta=0,\\[6.45831pt] \displaystyle\frac{r_{-}(a_{0}-r_{+})\exp\left(\frac{8\Gamma\Omega^{3}}{3}(r_{-}-r_{+})\hat{x}\right)-r_{+}(a_{0}-r_{-})}{(a_{0}-r_{+})\exp\left(\frac{8\Gamma\Omega^{3}}{3}(r_{-}-r_{+})\hat{x}\right)-(a_{0}-r_{-})},&\Delta<0.\end{cases}$ (25)
Because the relationship between $\Delta$ and $\delta$, we can rewrite the condition as function of $\delta$ as in Eq. 9.
We define $\hat{x}^{*}$ as the location where $U(\hat{x}^{*})=|v|$, or equivalently $a(\hat{x}^{*})\Omega=|v|$. Therefore:
$\hat{x}^{*}=\max\!\left\{\,a^{-1}\!\left(\frac{|v|}{\Omega}\right),\,0\,\right\},$ (26)
denotes the extent of region A.
B.2 Region B
In region B we enforce continuity of the amplitude at $\hat{x}^{*}$. Note that if the orbital velocity at the ice edge is lower that the drift only region B exists. Using the integral approximation, the differential and integral form of the governing equation are:
$\frac{da_{B}}{d\hat{x}}\;=\;-\left(\alpha+3\pi\Gamma\Omega^{2}|v|\right)a_{B}\;-\;\frac{2\pi\Gamma|v|^{3}}{a_{B}},$ (27)
$\int_{a_{A}(x^{*})}^{a_{B}}\frac{a^{\prime}}{\left(\alpha+3\pi\Gamma\Omega^{2}|v|\right)a^{\prime 2}+2\pi\Gamma|v|^{3}}\,da^{\prime}\;=\;x^{*}-\hat{x}$. (28)
The integrand yields:
$a_{B}(\hat{x})=\left[e^{-(6\pi\Omega^{2}|v|\Gamma+2\alpha)(\hat{x}-\hat{x}^{*})}\left(\frac{v^{2}}{\Omega^{2}}+\frac{2\pi|v|^{3}\Gamma}{3\pi\Omega^{2}|v|\Gamma+\alpha}\right)-\frac{2\pi|v|^{3}\Gamma}{3\pi\Omega^{2}|v|\Gamma+\alpha}\right]^{\tfrac{1}{2}},\qquad\hat{x}\geq\hat{x}^{*}.$ (29)